# Mapping 1,000+ Language Models via the Log-Likelihood Vector

### v1

@@ 7 Conclusion
We propose a method to compare autoregressive language models using log-likelihoods from a predefined text set. By treating these as model coordinates, we show that the squared Euclidean distance approximates KL divergence, enabling efficient large-scale comparisons. Experiments with over 1,000 models confirm its effectiveness in structuring model relationships and predicting benchmark task performance, while also validating the theoretical foundations.
@@
  • The results of the task-performance prediction in Section 5 should be interpreted conservatively. While we use a proper cross-validation setup to prevent data leakage (splitting the training set, validation set, and test set), if very similar models appear in both training and test sets (for example, models that share a base model and differ only slightly by fine-tuning), it might become easier to predict task performance. Also, we have not evaluated models that are not listed on the leaderboard.
@@ Acknowledgments
This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110, JST BOOST JPMJBS2407.
@@ Appendix A Related Work
In recent years, research on comparing large language models (LLMs) has gained attention. This section provides an overview of existing studies from three perspectives: model parameters, activations [^12] In general, “activations” refer to the intermediate outputs of Transformer models (e.g., the residual stream or neurons) Bereska and Gavves (2024)., and probability distributions.
@@
One approach to comparing LLMs is to analyze their parameters. Zhu et al. (2025) proposed a statistical framework for evaluating parameter similarity between different models and introduced a method for determining whether these models were trained independently. Additionally, Yadav et al. (2023b) focused on parameter changes due to task adaptation, specifically analyzing task vectors [^13] The difference in parameters before and after fine-tuning is referred to as a task vector, and arithmetic operations on these vectors are effective Ilharco et al. (2023).. They proposed a method to mitigate interference when integrating task vectors from different models. Specifically, by reducing redundant numerical components and adjusting for conflicting signs, their approach enables effective model merging.
@@
Comparison of probability distributions.
Several approaches compare LLMs using probability distributions. Lv et al. (2023) proposed a method for computing coefficients in parameter ensembling by providing the same input text to two models and comparing the softmax probability distributions at each token. Specifically, they used KL divergence and summed the results to derive appropriate coefficients. Furthermore, Yax et al. (2025) proposed a similarity metric based on the conditional probabilities of LLMs and introduced a method for calculating the phylogenetic distance between different models. Additionally, a method has been proposed for measuring differences in conditional probabilities based on prompts using KL divergence Melamed et al. (2024).
@@ Appendix I Table of Nearest Neighbor Models
bigcode/starcoder2-7b KL (/byte) codellama/CodeLlama-7b-Instruct-hf KL (/byte) bigcode/starcoder2-3b 769.780 (1.142) NousResearch/CodeLlama-7b-hf 48.192 (0.072) stabilityai/stable-code-3b 1606.483 (2.384) codellama/CodeLlama-7b-hf 48.192 (0.072) deepseek-ai/deepseek-coder-6.7b-base 1687.186 (2.503) codellama/CodeLlama-13b-Instruct-hf 138.606 (0.206) EleutherAI/llemma_7b 1892.812 (2.808) TheBloke/CodeLlama-13B-Instruct-fp16 138.627 (0.206) deepseek-ai/deepseek-coder-7b-base-v1.5 2031.482 (3.014) Nexusflow/NexusRaven-V2-13B 147.114 (0.218) deepseek-ai/DeepSeek-Coder-V2-Lite-Base 2042.158 (3.030) codellama/CodeLlama-13b-hf 158.786 (0.236) deepseek-ai/deepseek-coder-7b-instruct-v1.5 2045.330 (3.035) NousResearch/CodeLlama-13b-hf 158.949 (0.236) deepseek-ai/deepseek-coder-6.7b-instruct 2140.788 (3.176) OpenAssistant/codellama-13b-oasst-sft-v10 219.381 (0.326) deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct 2149.008 (3.189) HiTZ/GoLLIE-7B 225.804 (0.335) meta-math/MetaMath-Llemma-7B 2184.274 (3.241) WhiteRabbitNeo/WhiteRabbitNeo-13B-v1 260.047 (0.386) deepseek-ai/deepseek-coder-1.3b-base KL (/byte) deepseek-ai/deepseek-llm-7b-base KL (/byte) deepseek-ai/deepseek-coder-1.3b-instruct 410.897 (0.610) deepseek-ai/deepseek-moe-16b-base 130.715 (0.194) deepseek-ai/deepseek-coder-6.7b-instruct 512.997 (0.761) deepseek-ai/deepseek-llm-7b-chat 245.408 (0.364) deepseek-ai/deepseek-coder-6.7b-base 601.002 (0.892) deepseek-ai/deepseek-moe-16b-chat 247.889 (0.368) bigcode/starcoderbase-1b 798.322 (1.185) deepseek-ai/DeepSeek-V2-Lite 306.530 (0.455) bigcode/starcoderbase-7b 1249.899 (1.855) deepseek-ai/ESFT-vanilla-lite 307.487 (0.456) NTQAI/Nxcode-CQ-7B-orpo 1307.618 (1.940) deepseek-ai/DeepSeek-V2-Lite-Chat 585.090 (0.868) Qwen/CodeQwen1.5-7B-Chat 1317.166 (1.954) mistralai/Mistral-7B-Instruct-v0.1 913.708 (1.356) bigcode/starcoder2-3b 1858.329 (2.757) statking/zephyr-7b-sft-full-orpo 1089.127 (1.616) Salesforce/codegen-6B-multi 1974.095 (2.929) Severian/ANIMA-Phi-Neptune-Mistral-7B 1140.161 (1.692) google/codegemma-2b 2115.462 (3.139) sethuiyer/Medichat-Llama3-8B 1157.599 (1.718) EleutherAI/gpt-neo-1.3B KL (/byte) EleutherAI/pythia-12b KL (/byte) EleutherAI/gpt-neo-2.7B 218.723 (0.325) matsuo-lab/weblab-10b 139.270 (0.207) EleutherAI/pythia-1.4b 289.377 (0.429) h2oai/h2ogpt-oig-oasst1-256-6_9b 159.560 (0.237) EleutherAI/pythia-1b-deduped 413.204 (0.613) Salesforce/codegen-6B-nl 175.546 (0.260) HWERI/pythia-1.4b-deduped-sharegpt 421.551 (0.625) EleutherAI/gpt-j-6b 186.613 (0.277) beaugogh/pythia-1.4b-deduped-sharegpt 421.551 (0.625) TehVenom/Dolly_Malion-6b 231.701 (0.344) EleutherAI/pythia-1.4b-deduped 447.122 (0.663) TehVenom/PPO_Shygmalion-6b 233.159 (0.346) PygmalionAI/metharme-1.3b 448.970 (0.666) TehVenom/Dolly_Shygmalion-6b-Dev_V8P2 236.753 (0.351) RWKV/rwkv-raven-1b5 512.789 (0.761) TehVenom/PPO_Pygway-V8p4_Dev-6b 237.270 (0.352) databricks/dolly-v2-3b 525.574 (0.780) TehVenom/GPT-J-Pyg_PPO-6B-Dev-V8p4 245.015 (0.364) EleutherAI/pythia-2.8b-deduped 587.699 (0.872) TehVenom/PPO_Shygmalion-V8p4_Dev-6b 245.364 (0.364) facebook/opt-6.7b KL (/byte) google/codegemma-2b KL (/byte) KoboldAI/OPT-6.7B-Erebus 0.685 (0.001) deepseek-ai/deepseek-coder-1.3b-instruct 1654.767 (2.455) KoboldAI/OPT-6.7B-Nerybus-Mix 78.604 (0.117) bigcode/starcoderbase-1b 1812.768 (2.690) KoboldAI/OPT-6B-nerys-v2 124.520 (0.185) deepseek-ai/deepseek-coder-1.3b-base 2115.462 (3.139) facebook/opt-13b 196.634 (0.292) bigcode/gpt_bigcode-santacoder 2644.348 (3.924) KoboldAI/OPT-13B-Nerybus-Mix 230.046 (0.341) deepseek-ai/deepseek-coder-6.7b-instruct 2675.004 (3.969) KoboldAI/OPT-13B-Erebus 237.845 (0.353) Qwen/CodeQwen1.5-7B-Chat 2759.243 (4.094) KoboldAI/OPT-13B-Nerys-v2 273.410 (0.406) NTQAI/Nxcode-CQ-7B-orpo 2767.578 (4.106) facebook/opt-2.7b 327.851 (0.486) Salesforce/codegen-6B-multi 2859.701 (4.243) KoboldAI/OPT-2.7B-Nerybus-Mix 427.279 (0.634) bigcode/starcoderbase-7b 2990.326 (4.437) KoboldAI/OPT-2.7B-Erebus 446.712 (0.663) deepseek-ai/deepseek-coder-6.7b-base 3282.352 (4.870) google/gemma-7b KL (/byte) lmsys/vicuna-13b-v1.3 KL (/byte) SeaLLMs/SeaLLM-7B-v2.5 247.533 (0.367) TheBloke/stable-vicuna-13B-HF 195.152 (0.290) VAGOsolutions/SauerkrautLM-Gemma-7b 257.986 (0.383) Yhyu13/chimera-inst-chat-13b-hf 197.130 (0.292) lemon-mint/gemma-ko-7b-instruct-v0.62 344.582 (0.511) junelee/wizard-vicuna-13b 219.148 (0.325) google/gemma-2b 503.918 (0.748) TheBloke/wizard-vicuna-13B-HF 219.148 (0.325) google/codegemma-7b 596.383 (0.885) TheBloke/UltraLM-13B-fp16 219.689 (0.326) PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct 814.680 (1.209) NousResearch/Nous-Hermes-13b 221.324 (0.328) FlagAlpha/Llama3-Chinese-8B-Instruct 817.954 (1.214) project-baize/baize-v2-13b 233.813 (0.347) FairMind/Llama-3-8B-4bit-UltraChat-Ita 832.782 (1.236) TheBloke/guanaco-13B-HF 234.090 (0.347) migtissera/Llama-3-8B-Synthia-v3.5 842.861 (1.251) openaccess-ai-collective/minotaur-13b-fixed 235.470 (0.349) Orenguteng/Llama-3-8B-Lexi-Uncensored 844.748 (1.253) openaccess-ai-collective/wizard-mega-13b 238.073 (0.353) medalpaca/medalpaca-7b KL (/byte) meta-llama/Llama-2-13b-hf KL (/byte) TheBloke/guanaco-7B-HF 285.806 (0.424) TaylorAI/Flash-Llama-13B 0.000 (0.000) eachadea/vicuna-7b-1.1 336.518 (0.499) TheBloke/Llama-2-13B-fp16 0.002 (0.000) TehVenom/Pygmalion-Vicuna-1.1-7b 369.003 (0.548) StudentLLM/Alpagasus-2-13b-QLoRA-merged 4.502 (0.007) lmsys/vicuna-7b-v1.3 374.469 (0.556) CHIH-HUNG/llama-2-13b-FINETUNE2_TEST_2.2w 7.610 (0.011) jphme/orca_mini_v2_ger_7b 383.382 (0.569) garage-bAInd/Platypus2-13B 10.373 (0.015) ajibawa-2023/Uncensored-Jordan-7B 384.396 (0.570) CHIH-HUNG/llama-2-13b-dolphin_5w 14.386 (0.021) bofenghuang/vigogne-7b-instruct 390.797 (0.580) CHIH-HUNG/llama-2-13b-OpenOrca_5w 14.982 (0.022) TheBloke/airoboros-7b-gpt4-fp16 391.991 (0.582) CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r16-gate_up_down-test1 17.256 (0.026) TheBloke/tulu-7B-fp16 423.476 (0.628) CHIH-HUNG/llama-2-13b-FINETUNE4_addto15k_4.5w-r16-gate_up_down 17.405 (0.026) TehVenom/Pygmalion_AlpacaLora-7b 431.540 (0.640) CHIH-HUNG/llama-2-13b-FINETUNE4_compare15k_4.5w-r16-gate_up_down 19.427 (0.029) meta-llama/Llama-2-7b-hf KL (/byte) meta-llama/Meta-Llama-3-8B KL (/byte) meta-llama/Llama-2-7b-hf 0.000 (0.000) Undi95/Meta-Llama-3-8B-hf 0.003 (0.000) TheTravellingEngineer/llama2-7b-chat-hf-v4 0.000 (0.000) dfurman/Llama-3-8B-Orpo-v0.1 6.994 (0.010) ibranze/araproje-llama2-7b-hf 0.000 (0.000) migtissera/Tess-2.0-Llama-3-8B 28.829 (0.043) yeen214/test_llama2_7b 0.000 (0.000) freewheelin/free-llama3-dpo-v0.2 67.783 (0.101) TheTravellingEngineer/llama2-7b-chat-hf-v2 0.000 (0.000) jondurbin/bagel-8b-v1.0 110.365 (0.164) NewstaR/Starlight-7B 0.003 (0.000) migtissera/Llama-3-8B-Synthia-v3.5 128.141 (0.190) Delcos/Mistral-Pygmalion-7b 14.809 (0.022) nvidia/Llama3-ChatQA-1.5-8B 128.776 (0.191) elliotthwang/elliott_Llama-2-7b-hf 16.612 (0.025) ruslanmv/Medical-Llama3-8B 138.566 (0.206) garage-bAInd/Platypus2-7B 22.811 (0.034) FairMind/Llama-3-8B-4bit-UltraChat-Ita 199.748 (0.296) Lazycuber/L2-7b-Base-Guanaco-Uncensored 23.345 (0.035) NousResearch/Hermes-2-Theta-Llama-3-8B 222.357 (0.330) meta-math/MetaMath-Mistral-7B KL (/byte) mistralai/Mistral-7B-v0.3 KL (/byte) Weyaxi/MetaMath-NeuralHermes-2.5-Mistral-7B-Linear 43.095 (0.064) MaziyarPanahi/Mistral-7B-v0.3 0.000 (0.000) Weyaxi/MetaMath-Tulpar-7b-v2-Slerp 81.882 (0.121) mistral-community/Mistral-7B-v0.2 7.221 (0.011) Weyaxi/MetaMath-OpenHermes-2.5-neural-chat-v3-3-Slerp 101.097 (0.150) unsloth/mistral-7b-v0.2 7.222 (0.011) Q-bert/Bumblebee-7B 106.376 (0.158) mistralai/Mistral-7B-v0.1 22.038 (0.033) OpenPipe/mistral-ft-optimized-1227 110.089 (0.163) Cartinoe5930/Llama2_init_Mistral 29.819 (0.044) Toten5/Marcoroni-neural-chat-7B-v2 114.004 (0.169) Locutusque/Hercules-3.1-Mistral-7B 32.099 (0.048) ignos/Mistral-T5-7B-v1 114.320 (0.170) migtissera/Synthia-7B-v3.0 33.418 (0.050) Weyaxi/MetaMath-Chupacabra-7B-v2.01-Slerp 114.414 (0.170) uukuguy/speechless-zephyr-code-functionary-7b 35.723 (0.053) Q-bert/Optimus-7B 117.829 (0.175) uukuguy/zephyr-7b-alpha-dare-0.85 35.745 (0.053) Weyaxi/MetaMath-NeuralHermes-2.5-Mistral-7B-Ties 127.457 (0.189) crumb/apricot-wildflower-20 40.907 (0.061)
@@
Table 5: Top 10 nearest neighbors among the 1,018 language models for the models labeled in the top panel of Fig. 1. The KL divergence is computed using formula (3) in Section 2. The values in parentheses represent the KL divergence measured in bits per byte of text (Section 3.5), obtained by multiplying the KL divergence by 0.001484.
@@
Figure 18: Scatter plots of predicted scores and benchmark scores for six benchmark tasks. Additionally, results for predicting 6-TaskMean (identical to Fig. 8) and the mean log-likelihood are also shown. Each point is color-coded by the mean log-likelihood, with higher mean log-likelihood values generally corresponding to higher task scores. For better visualization, the color bar range is clipped to the 10th–100th percentile.
@@
For each composed model $p_{\alpha,\beta}$ with weight $W_{\alpha,\beta}$, we computed the model coordinates $\bm{q}_{\alpha,\beta}$ following the method described in Section 2. The tokenizer of the base model was used to ensure consistency. The text data consists of 1,000 randomly sampled sentences from the dataset of 10,000 sentences created in Section 3.1.
@@
We set the base model $p_{0}$ as Llama-2-7b-hf. Among the fine-tuned models based on Llama-2-7b-hf, we selected the two most downloaded models: vicuna-7b-v1.5 and Llama-2-7b-chat-hf, denoted as $p_{1}$ and $p_{2}$, respectively.
@@
Figure 10 shows the linearly merged models, visualized in both weight space and log-likelihood space. The corners of the grid are labeled with their corresponding $(\alpha,\beta)$ values.

### v2

@@ 7 Conclusion
We proposed a method to compare autoregressive language models using log-likelihoods on a predefined text set. By interpreting these as model coordinates, we showed that the squared Euclidean distance approximates KL divergence. Experiments on over 1,000 models confirmed the method’s effectiveness for analyzing model relationships, predicting benchmark performance, and validating the theory.
@@
  • The results of the task-performance prediction in Section 5 should be interpreted conservatively. We employed a proper cross-validation, splitting the models based on their model types so that the training, validation, and test sets were disjoint, thereby limiting data leakage. However, if nearly identical models appear in both the training and test partitions (for example, models labeled with different types but built on the same base model and modified only slightly by fine-tuning), prediction becomes artificially easier. As shown in Tables 7 and 8 in Appendix K.3, random splits that permit such leakage yield higher predictive performance than splits that strictly follow model-type groupings. Finally, note that models not listed on the leaderboard were excluded from our evaluation.
@@ Acknowledgments
This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST BOOST JPMJBS2407.
@@ Appendix A Related Work
In recent years, research on comparing large language models (LLMs) has gained attention. This section provides an overview of existing studies from three perspectives: model parameters, activations [^13] In general, “activations” refer to the intermediate outputs of Transformer models (e.g., the residual stream or neurons) Bereska and Gavves (2024)., and model outputs.
@@
One approach to comparing LLMs is to analyze their parameters. Zhu et al. (2025) proposed a statistical framework for evaluating parameter similarity between different models and introduced a method for determining whether these models were trained independently. Horwitz et al. (2025) compare weights to complement undocumented model relationships on Hugging Face. Additionally, Yadav et al. (2023b) focused on parameter changes due to task adaptation, specifically analyzing task vectors [^14] The difference in parameters before and after fine-tuning is referred to as a task vector, and arithmetic operations on these vectors are effective Ilharco et al. (2023).. They proposed a method to mitigate interference when integrating task vectors from different models. Specifically, by reducing redundant numerical components and adjusting for conflicting signs, their approach enables effective model merging.
@@
Comparison of model outputs.
Several approaches compare LLMs based on their outputs. Lv et al. (2023) proposed a method for computing coefficients in parameter ensembling by providing the same input text to two models and comparing the softmax probability distributions at each token. Specifically, they used KL divergence and summed the results to derive appropriate coefficients. Furthermore, Yax et al. (2025) proposed a similarity metric based on the conditional probabilities of LLMs and introduced a method for calculating the phylogenetic distance between different models. Zhuang et al. (2025) propose a framework for vector representations of language models based on their performance on downstream tasks. Additionally, a method has been proposed for measuring differences in conditional probabilities based on prompts using KL divergence Melamed et al. (2024).
@@ Appendix I Table of Nearest Neighbor Models
bigcode/starcoder2-7b KL [bpb] codellama/CodeLlama-7b-Instruct-hf KL [bpb] bigcode/starcoder2-3b 1.14 codellama/CodeLlama-7b-hf 0.0715 stabilityai/stable-code-3b 2.38 NousResearch/CodeLlama-7b-hf 0.0715 deepseek-ai/deepseek-coder-6.7b-base 2.50 codellama/CodeLlama-13b-Instruct-hf 0.206 EleutherAI/llemma_7b 2.81 TheBloke/CodeLlama-13B-Instruct-fp16 0.206 deepseek-ai/deepseek-coder-7b-base-v1.5 3.01 Nexusflow/NexusRaven-V2-13B 0.218 deepseek-ai/DeepSeek-Coder-V2-Lite-Base 3.03 codellama/CodeLlama-13b-hf 0.236 deepseek-ai/deepseek-coder-7b-instruct-v1.5 3.03 NousResearch/CodeLlama-13b-hf 0.236 deepseek-ai/deepseek-coder-6.7b-instruct 3.18 OpenAssistant/codellama-13b-oasst-sft-v10 0.326 deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct 3.19 HiTZ/GoLLIE-7B 0.335 meta-math/MetaMath-Llemma-7B 3.24 WhiteRabbitNeo/WhiteRabbitNeo-13B-v1 0.386 deepseek-ai/deepseek-coder-1.3b-base KL [bpb] deepseek-ai/deepseek-llm-7b-base KL [bpb] deepseek-ai/deepseek-coder-1.3b-instruct 0.610 deepseek-ai/deepseek-moe-16b-base 0.194 deepseek-ai/deepseek-coder-6.7b-instruct 0.761 deepseek-ai/deepseek-llm-7b-chat 0.364 deepseek-ai/deepseek-coder-6.7b-base 0.892 deepseek-ai/deepseek-moe-16b-chat 0.368 bigcode/starcoderbase-1b 1.185 deepseek-ai/DeepSeek-V2-Lite 0.455 bigcode/starcoderbase-7b 1.855 deepseek-ai/ESFT-vanilla-lite 0.456 NTQAI/Nxcode-CQ-7B-orpo 1.940 deepseek-ai/DeepSeek-V2-Lite-Chat 0.868 Qwen/CodeQwen1.5-7B-Chat 1.954 mistralai/Mistral-7B-Instruct-v0.1 1.356 bigcode/starcoder2-3b 2.757 statking/zephyr-7b-sft-full-orpo 1.616 Salesforce/codegen-6B-multi 2.929 Severian/ANIMA-Phi-Neptune-Mistral-7B 1.692 google/codegemma-2b 3.139 sethuiyer/Medichat-Llama3-8B 1.718 EleutherAI/gpt-neo-1.3B KL [bpb] EleutherAI/pythia-12b KL [bpb] EleutherAI/gpt-neo-2.7B 0.325 matsuo-lab/weblab-10b 0.207 EleutherAI/pythia-1.4b 0.429 h2oai/h2ogpt-oig-oasst1-256-6_9b 0.237 EleutherAI/pythia-1b-deduped 0.613 Salesforce/codegen-6B-nl 0.260 HWERI/pythia-1.4b-deduped-sharegpt 0.625 EleutherAI/gpt-j-6b 0.277 beaugogh/pythia-1.4b-deduped-sharegpt 0.625 TehVenom/Dolly_Malion-6b 0.344 EleutherAI/pythia-1.4b-deduped 0.663 TehVenom/PPO_Shygmalion-6b 0.346 PygmalionAI/metharme-1.3b 0.666 TehVenom/Dolly_Shygmalion-6b-Dev_V8P2 0.351 RWKV/rwkv-raven-1b5 0.761 TehVenom/PPO_Pygway-V8p4_Dev-6b 0.352 databricks/dolly-v2-3b 0.780 TehVenom/GPT-J-Pyg_PPO-6B-Dev-V8p4 0.364 EleutherAI/pythia-2.8b-deduped 0.872 TehVenom/PPO_Shygmalion-V8p4_Dev-6b 0.364 facebook/opt-6.7b KL [bpb] google/codegemma-2b KL [bpb] KoboldAI/OPT-6.7B-Erebus 0.00102 deepseek-ai/deepseek-coder-1.3b-instruct 2.46 KoboldAI/OPT-6.7B-Nerybus-Mix 0.117 bigcode/starcoderbase-1b 2.69 KoboldAI/OPT-6B-nerys-v2 0.185 deepseek-ai/deepseek-coder-1.3b-base 3.14 facebook/opt-13b 0.292 bigcode/gpt_bigcode-santacoder 3.92 KoboldAI/OPT-13B-Nerybus-Mix 0.341 deepseek-ai/deepseek-coder-6.7b-instruct 3.97 KoboldAI/OPT-13B-Erebus 0.353 Qwen/CodeQwen1.5-7B-Chat 4.09 KoboldAI/OPT-13B-Nerys-v2 0.406 NTQAI/Nxcode-CQ-7B-orpo 4.11 facebook/opt-2.7b 0.486 Salesforce/codegen-6B-multi 4.24 KoboldAI/OPT-2.7B-Nerybus-Mix 0.634 bigcode/starcoderbase-7b 4.44 KoboldAI/OPT-2.7B-Erebus 0.663 deepseek-ai/deepseek-coder-6.7b-base 4.87 google/gemma-7b KL [bpb] lmsys/vicuna-13b-v1.3 KL [bpb] SeaLLMs/SeaLLM-7B-v2.5 0.367 TheBloke/stable-vicuna-13B-HF 0.290 VAGOsolutions/SauerkrautLM-Gemma-7b 0.383 Yhyu13/chimera-inst-chat-13b-hf 0.292 lemon-mint/gemma-ko-7b-instruct-v0.62 0.511 junelee/wizard-vicuna-13b 0.325 google/gemma-2b 0.748 TheBloke/wizard-vicuna-13B-HF 0.325 google/codegemma-7b 0.885 TheBloke/UltraLM-13B-fp16 0.326 PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct 1.209 NousResearch/Nous-Hermes-13b 0.328 FlagAlpha/Llama3-Chinese-8B-Instruct 1.214 project-baize/baize-v2-13b 0.347 FairMind/Llama-3-8B-4bit-UltraChat-Ita 1.236 TheBloke/guanaco-13B-HF 0.347 migtissera/Llama-3-8B-Synthia-v3.5 1.251 openaccess-ai-collective/minotaur-13b-fixed 0.349 Orenguteng/Llama-3-8B-Lexi-Uncensored 1.253 openaccess-ai-collective/wizard-mega-13b 0.353 medalpaca/medalpaca-7b KL [bpb] meta-llama/Llama-2-13b-hf KL [bpb] TheBloke/guanaco-7B-HF 0.424 TaylorAI/Flash-Llama-13B 1.82e-16 eachadea/vicuna-7b-1.1 0.499 TheBloke/Llama-2-13B-fp16 3.6e-06 TehVenom/Pygmalion-Vicuna-1.1-7b 0.548 StudentLLM/Alpagasus-2-13b-QLoRA-merged 0.00668 lmsys/vicuna-7b-v1.3 0.556 CHIH-HUNG/llama-2-13b-FINETUNE2_TEST_2.2w 0.0113 jphme/orca_mini_v2_ger_7b 0.569 garage-bAInd/Platypus2-13B 0.0154 ajibawa-2023/Uncensored-Jordan-7B 0.570 CHIH-HUNG/llama-2-13b-dolphin_5w 0.0213 bofenghuang/vigogne-7b-instruct 0.580 CHIH-HUNG/llama-2-13b-OpenOrca_5w 0.0222 TheBloke/airoboros-7b-gpt4-fp16 0.582 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r16-gate_up_down-test1 0.0256 TheBloke/tulu-7B-fp16 0.628 CHIH-HUNG/llama-2-13b-FINETUNE4_addto15k_4.5w-r16-gate_up_down 0.0258 TehVenom/Pygmalion_AlpacaLora-7b 0.640 CHIH-HUNG/llama-2-13b-FINETUNE4_compare15k_4.5w-r16-gate_up_down 0.0288 meta-llama/Llama-2-7b-hf KL [bpb] meta-llama/Meta-Llama-3-8B KL [bpb] ibranze/araproje-llama2-7b-hf 0 Undi95/Meta-Llama-3-8B-hf 4.56e-06 TheTravellingEngineer/llama2-7b-chat-hf-v4 0 dfurman/Llama-3-8B-Orpo-v0.1 0.0104 TheTravellingEngineer/llama2-7b-chat-hf-v2 0 migtissera/Tess-2.0-Llama-3-8B 0.0428 TaylorAI/Flash-Llama-7B 0 freewheelin/free-llama3-dpo-v0.2 0.101 yeen214/test_llama2_7b 0 jondurbin/bagel-8b-v1.0 0.164 NewstaR/Starlight-7B 4.57e-06 migtissera/Llama-3-8B-Synthia-v3.5 0.190 Delcos/Mistral-Pygmalion-7b 0.0220 nvidia/Llama3-ChatQA-1.5-8B 0.191 elliotthwang/elliott_Llama-2-7b-hf 0.0246 ruslanmv/Medical-Llama3-8B 0.206 garage-bAInd/Platypus2-7B 0.0338 FairMind/Llama-3-8B-4bit-UltraChat-Ita 0.296 Lazycuber/L2-7b-Base-Guanaco-Uncensored 0.0346 NousResearch/Hermes-2-Theta-Llama-3-8B 0.330 meta-math/MetaMath-Mistral-7B KL [bpb] mistralai/Mistral-7B-v0.3 KL [bpb] Weyaxi/MetaMath-NeuralHermes-2.5-Mistral-7B-Linear 0.0639 MaziyarPanahi/Mistral-7B-v0.3 3.64e-16 Weyaxi/MetaMath-Tulpar-7b-v2-Slerp 0.121 mistral-community/Mistral-7B-v0.2 0.0107 Weyaxi/MetaMath-OpenHermes-2.5-neural-chat-v3-3-Slerp 0.150 unsloth/mistral-7b-v0.2 0.0107 Q-bert/Bumblebee-7B 0.158 mistralai/Mistral-7B-v0.1 0.0327 OpenPipe/mistral-ft-optimized-1227 0.163 Cartinoe5930/Llama2_init_Mistral 0.0442 Toten5/Marcoroni-neural-chat-7B-v2 0.169 Locutusque/Hercules-3.1-Mistral-7B 0.0476 ignos/Mistral-T5-7B-v1 0.170 migtissera/Synthia-7B-v3.0 0.0496 Weyaxi/MetaMath-Chupacabra-7B-v2.01-Slerp 0.170 uukuguy/speechless-zephyr-code-functionary-7b 0.0530 Q-bert/Optimus-7B 0.175 uukuguy/zephyr-7b-alpha-dare-0.85 0.0530 Weyaxi/MetaMath-NeuralHermes-2.5-Mistral-7B-Ties 0.189 crumb/apricot-wildflower-20 0.0607
@@
Table 6: Top 10 nearest neighbors among the 1,018 language models for each model labeled in the top panel of Fig. 1. The values indicate the KL divergence measured in bits per byte (bpb), as defined in Section 3.5. These are computed using formula (3) in Section 2, by multiplying the original KL divergence by 0.001484.
@@
Figure 18: Scatter plots of predicted scores versus benchmark scores for test sets across the six benchmark tasks (the dashed line indicates the identity line). Additionally, results for predicting 6-TaskMean (identical to Fig. 8) and the mean log-likelihood are also shown. Each point is color-coded by the mean log-likelihood, with higher mean log-likelihood values generally corresponding to higher task scores. For better visualization, the color bar range is clipped to the 10th–100th percentile.
@@
For each composed model $p_{\alpha,\beta}$ with weight $W_{\alpha,\beta}$, we computed the model coordinates $\bm{q}_{\alpha,\beta}$ following the method described in Section 2. The tokenizer of the base model was used to ensure consistency. The text data consists of all 10,000 texts prepared in Section 3.1. Additionally, the deterministic algorithms option202020torch.use_deterministic_algorithms(True) was enabled in the implementation to ensure reproducibility.
@@
[…]
@@
Figure 19: Visualization of 36 language models obtained by linearly interpolating pretrained model weights based on mistralai/Mistral-7B-v0.1. Each point is color-coded according to its mean log-likelihood. (Left) Models in the weight parameter space. (Right) Models in the log-likelihood space, represented by the $\bm{q}$-coordinate system.
Figure 20: Scatter plots comparing the actual and predicted mean log-likelihoods for 36 interpolated models derived from Llama-2-7b-hf (left) and Mistral-7B-v0.1 (right). Each point corresponds to a unique combination of merge ratios $(\alpha,\beta)$ and is color-coded accordingly. The dashed line indicates ideal prediction. Strong correlations are observed for both model sets: Llama-2 shows $r=0.969$, $\rho=0.996$, and Mistral shows $r=0.973$, $\rho=0.975$. A 2D color map of $(\alpha,\beta)$ values is shown as an inset.
Figures 10 and 19 show the linearly merged models, visualized in both weight space and log-likelihood space. The corners of the grid are labeled with their corresponding $(\alpha,\beta)$ values.
@@
Estimating the mean log-likelihood of interpolated models.
To assess the potential of predicting task performance from log-likelihood vectors without explicitly computing them for every interpolated model, we evaluated whether the mean log-likelihoods can be estimated by linearly interpolating those of the base models. Let $\ell_{\alpha,\beta}(x)$ denote the log-likelihood of input $x$ computed by the merged model $p_{\alpha,\beta}$. Let $\ell_{0}$, $\ell_{1}$, and $\ell_{2}$ be the log-likelihoods from the base and fine-tuned models $p_{0}$, $p_{1}$, and $p_{2}$, respectively. We define the linearly interpolated log-likelihood as:
^$\hat{\ell}_{\alpha,\beta}=\ell_{0}+\alpha(\ell_{1}-\ell_{0})+\beta(\ell_{2}-\ell_{0}).$ (43)
Figure 20 shows a strong correlation between the predicted and actual mean log-likelihoods. This result suggests that log-likelihood vectors can be approximated by linear interpolation, enabling efficient prediction of model performance without computing log-likelihoods for every model.